### v2

# A $\mathbb{CP}^{2}$ SMEFT

Abstract
An extension of the Standard Model is proposed, where the Higgs field is valued in the complex projective plane $\mathbb{CP}^{2}$, rather than $\mathbb{C}^{2}$. Its geometry is consistent with $U(2)\simeq(SU(2)\times U(1))/\mathbb{Z}_{2}$ electroweak gauge symmetry. The leading terms in the Lagrangian, beyond those of the Standard Model, are much more tightly constrained than in general SMEFTs, and the custodial $SO(4)$ symmetry of the Higgs sector is mildly broken. The predicted tree-level deviations from Standard Model phenomenology depend on a single large mass parameter $M$. Current experimental data imply that $M$ is a few TeV or larger.
1 Introduction
We propose here an extension of the Standard Model (SM) [1] in the spirit of a SMEFT (Standard Model Effective Field Theory) [2]. Its perturbative field expansion has terms of all orders, but the dominant non-renormalisable terms are of mass dimension 6. The key idea is to exploit the geometry of the complex projective plane $\mathbb{CP}^{2}$ as the manifold where the Higgs field takes its values. The Riemannian metric on $\mathbb{CP}^{2}$ is assumed to be the Fubini–Study metric, the standard maximally symmetric metric on $\mathbb{CP}^{2}$. In our model, its isometry group $SU(3)$ is present in the background, but explicitly broken by choosing a point on $\mathbb{CP}^{2}$ and retaining only the $U(2)$ isotropy group of this point as the gauge symmetry of the Lagrangian. In other words, our model is a $\mathbb{CP}^{2}$ sigma-model supplemented by a Higgs potential, where the $U(2)$ action is gauged.
Locally, near the chosen point of $\mathbb{CP}^{2}$, the geometry resembles that of flat $\mathbb{C}^{2}$ near the origin, and there is one complex-doublet Higgs field. Our $\mathbb{CP}^{2}$ model is therefore close to the Standard Model, provided the geometric scale factor of the $\mathbb{CP}^{2}$ manifold is large and the Higgs potential is chosen so that the Higgs vacuum manifold is small on this scale. In units of the Higgs vacuum expectation value (vev), the model therefore has a single large mass parameter $M$ for Beyond the Standard Model (BSM) physics.
In the electroweak sector of the Standard Model, including its gauge fields and fermions, the symmetry group is $U(2)\simeq(SU(2)\times U(1))/\mathbb{Z}_{2}$. However, the usual Higgs potential has a larger, custodial $SO(4)$ symmetry, and the Higgs vacuum manifold is a round 3-sphere. This custodial symmetry is not exploited here, so our model has a different geometric structure than models where it is retained [3, 4]. However, the generic orbits of $U(2)$ on $\mathbb{CP}^{2}$ are still three-dimensional; they are a family of squashed 3-spheres. One orbit is simply a point, and in the neighbourhood of this, the orbits are very close to being round 3-spheres. Therefore the custodial symmetry is approximately retained, and its important consequence for the ratio of the $W$-boson and $Z$-boson masses remains approximately valid. It was observed in ref.[3] that $SO(4)$ symmetry, broken to $U(2)$, leads to squashed 3-spheres, but the connection to $\mathbb{CP}^{2}$ geometry was not mentioned.
It is well known that there is no inconsistency at present in assuming that the SM gauge group is $(SU(3)\times SU(2)\times U(1))/\mathbb{Z}_{6}$. The quantum numbers of the Higgs field and the known fermion multiplets (quarks and leptons) lie in representations that are consistent with the $\mathbb{Z}_{6}$ quotient, and there are no local or global anomalies [5]. In the analysis that follows, we will ignore the QCD interactions of quarks and gluons. These sectors of the Standard Model Lagrangian are unaffected by changes to the Higgs field geometry. We will discuss only the electroweak sector, including the Higgs field, and the quark and lepton Yukawa couplings to the Higgs field. The relevant gauge group is therefore the electroweak $U(2)$ subgroup of the SM gauge group, mentioned above. We will also ignore fermion generation mixing, and simplify the neutrino physics to the version that was considered in the 1960s and ’70s, where there is just a left-handed neutrino field; our model has no novel insights concerning these challenging topics.
The Lagrangian has terms involving: (i) Gauge-covariant derivatives of the Higgs field, (ii) a Higgs potential, (iii) Gauge field tensors, (iv) Dirac (kinetic) terms for fermions, (v) Yukawa terms coupling the Higgs field and fermions. We will discuss these in turn, and then explore some phenomenological consequences, in particular, for the $W$-boson to $Z$-boson mass ratio, and for the Yukawa couplings of the Higgs boson. These will lead to constraints on the BSM mass parameter $M$.
Our presentation begins with a review of the Fubini–Study metric on $\mathbb{CP}^{2}$.
2 $\mathbb{CP}^{2}$ Geometry
The complex projective plane $\mathbb{CP}^{2}$ is obtained from flat complex 3-space, $\mathbb{C}^{3}$, by quotienting out the action of scaling by a complex number. Geometrically, $\mathbb{CP}^{2}$ is the manifold of complex lines [^2] We use the mathematician’s notion of complex line $\mathbb{C}$ to denote what is more commonly called the complex plane. All lines considered here are lines through the origin of $\mathbb{C}^{3}$. in $\mathbb{C}^{3}$.
We denote a non-zero point in $\mathbb{C}^{3}$ by the (column) vector $Z=(Z_{0},Z_{1},Z_{2})$ and the scaling action is $Z\to\kappa Z$ with $\kappa\in\mathbb{C}-\{0\}$. It is convenient to use the inhomogeneous coordinates $Z=(1,z_{1},z_{2})$ on $\mathbb{CP}^{2}$. However, this misses the part of $\mathbb{CP}^{2}$ “at infinity”. The missing part is the manifold of points $Z=(0,Z_{1},Z_{2})$ modulo scaling by $\kappa$, which is a copy of the complex projective line $\mathbb{CP}^{1}$, and geometrically a 2-sphere. Most of our calculations will be in the neighbourhood of $Z=(1,0,0)$ so the use of inhomogeneous coordinates causes no difficulty.
The hermitian conjugate (row) vector $\bar{Z}=(Z_{0}^{*},Z_{1}^{*},Z_{2}^{*})$ will play an important role in what follows. In inhomogeneous coordinates, the hermitian conjugate of $Z$ is $\bar{Z}=(1,z_{1}^{*},z_{2}^{*})$.
$\mathbb{C}^{3}$ has the Euclidean metric $ds^{2}=d\bar{Z}dZ$, whose isometry group, fixing the origin, is $SO(6)$. The natural metric on $\mathbb{CP}^{2}$ is derived from this by removing from $dZ$ its projection parallel to $Z$, and then normalising so that rescaling by $\kappa$ has no effect. The result is the Fubini–Study metric [6]
$ds^{2}=\frac{(\bar{Z}Z)d\bar{Z}dZ-(d\bar{Z}Z)(\bar{Z}dZ)}{(\bar{Z}Z)^{2}}\,.$ (1)
This is in dimensionless form; a scale factor will be introduced with the Higgs field. Note that if $dZ=\mu Z$ with $\mu\in\mathbb{C}$, i.e. $dZ$ parallel to $Z$, then $ds^{2}=0$.
The Fubini–Study metric is Kähler and of constant positive curvature. Because of the projection mentioned above, it depends in an essential way on the complex character of the coordinates. The isometry group is therefore not $SO(6)$ but its subgroup $SU(3)$. In homogeneous coordinates, the symmetry group action is simply $Z\to UZ$, with $U\in SU(3)$. $SU(3)$ acts transitively on $\mathbb{CP}^{2}$, and the isotropy group of any given point is $U(2)$. As a manifold, therefore, $\mathbb{CP}^{2}=SU(3)/U(2)$. We can check the isotropy group for the point $Z=(1,0,0)$, as follows. The matrices $U\in SU(3)$ that preserve $Z$ are of the form
$U=\begin{pmatrix}e^{-2i\gamma}&0\\ 0&e^{i\gamma}u\end{pmatrix}\,,$ (2)
with $u\in SU(2)$. Naively, this sends $(1,0,0)$ to $(e^{-2i\gamma},0,0)$, but this is equivalent to $(1,0,0)$ in the projective space. $e^{i\gamma}u$ is a unitary $2\times 2$ matrix, and apparently an element of $SU(2)\times U(1)$, but it is unchanged if $\gamma\to\gamma+\pi$ and $u\to-u$, so the isotropy group is indeed $U(2)$.
In our model, the point $(1,0,0)\in\mathbb{CP}^{2}$ plays a preferred role, and its isotropy group $U(2)$ will be identified with the electroweak gauge group. The generic orbits of $U(2)$ on $\mathbb{CP}^{2}$ are 3-spheres, as will be clarified in the next section.
Later, we will be using the projections of a generic vector $\Psi\in\mathbb{C}^{3}$, defined at location $Z\in\mathbb{C}^{3}$, on to the line through $Z$ and on to its complement, which is a copy of the tangent space to $\mathbb{CP}^{2}$. These projections are respectively
$\Psi\to P\Psi=\frac{Z(\bar{Z}\Psi)}{\bar{Z}Z}$ (3)
and
$\Psi\to(1-P)\Psi=\frac{(\bar{Z}Z)\Psi-Z(\bar{Z}\Psi)}{\bar{Z}Z}\,.$ (4)
Because these projections are unaffected by rescaling $Z$ by $\kappa$, they are defined at location $Z$ on $\mathbb{CP}^{2}$. More precisely, over $\mathbb{CP}^{2}$ there is its 1-dimensional tautological bundle, whose fibre at a point $Z\in\mathbb{CP}^{2}$ is the line in $\mathbb{C}^{3}$ through $Z$, and its 2-dimensional tangent bundle. The 3-dimensional, trivial $\mathbb{C}^{3}$ bundle over $\mathbb{CP}^{2}$ is the direct sum of these, so an element $\Psi$ in the trivial $\mathbb{C}^{3}$ bundle is the sum of its projections on to the tautological and tangent bundles.
Note that $\bar{Z}\Psi$ and $\bar{Z}Z$ are $SU(3)$-invariant. Therefore, both $P\Psi$ and $(1-P)\Psi$ transform in the same way as $\Psi$ under $SU(3)$. $P$ and $1-P$ themselves transform by conjugation under $SU(3)$. The projector $1-P$ will be important when we discuss Yukawa terms.
3 Squashed 3-Spheres
Let $z=(z_{1},z_{2})$ be standard coordinates on $\mathbb{C}^{2}$, and $ds^{2}=d\bar{z}dz$ the Euclidean metric on $\mathbb{C}^{2}$. Useful polar coordinates are
$z_{1}=r\cos\frac{1}{2}\theta\,e^{i(\chi-\varphi)/2}\,,\quad z_{2}=r\sin\frac{1}{2}\theta\,e^{i(\chi+\varphi)/2}\,,$ (5)
with ranges $0\leq r$, $0\leq\theta\leq\pi$, $0\leq\varphi<2\pi$ and $0\leq\chi<4\pi$. The submanifold of $\mathbb{C}^{2}$ at fixed $r>0$ is a round 3-sphere of radius $r$, which in turn is a circle bundle (a Hopf bundle) over a round 2-sphere, with $z_{2}/z_{1}=\tan\frac{1}{2}\theta\,e^{i\varphi}$ the stereographic coordinate on the 2-sphere, and $\chi$ the circle coordinate. In polars, the metric $ds^{2}=d\bar{z}dz$ becomes
$ds^{2}=dr^{2}+\frac{1}{4}r^{2}\left[d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}+(d\chi-\cos\theta\,d\varphi)^{2}\right],$ (6)
whose angular part is the metric of the round 3-sphere of radius $r$. Each of these 3-spheres is a $U(2)$ orbit. and there is a unique $U(2)$ fixed point at $r=0$. Note that the ratio of the circle radius to the 2-sphere radius is a constant, independent of $r$.
Let us now compare the Fubini–Study metric on $\mathbb{CP}^{2}$. Using the inhomogeneous coordinates $Z=(1,z_{1},z_{2})$, and hence $dZ=(0,dz_{1},dz_{2})$, the metric (1) becomes
$ds^{2}=\frac{(1+\bar{z}z)d\bar{z}dz-(d\bar{z}z)(\bar{z}dz)}{(1+\bar{z}z)^{2}}\,,$ (7)
and using the polar coordinates (5),
$ds^{2}=\frac{dr^{2}}{(1+r^{2})^{2}}+\frac{1}{4}\frac{r^{2}}{(1+r^{2})^{2}}\left[(1+r^{2})(d\theta^{2}+\sin^{2}\theta\,d\varphi^{2})+(d\chi-\cos\theta\,d\varphi)^{2}\right].$ (8)
Here, the angular part is the metric of a squashed 3-sphere, a Hopf circle bundle over a round 2-sphere, where the ratio of the circle radius to the 2-sphere radius is $1/\sqrt{1+r^{2}}$, i.e. dependent on $r$. For us, the important observation is that for small positive $r$, the radius ratio is close to unity, so the squashed 3-sphere is approximately round. The consequence for our model is that the custodial $SO(4)$ symmetry – the symmetry of a round 3-sphere – will be approximately realised by the Higgs potential.
As $r\to\infty$, the radius ratio approaches zero, compatibly with the completion of $\mathbb{CP}^{2}$ at infinity being simply a round 2-sphere, which is a special 2-dimensional orbit of $U(2)$. It has radius $\frac{1}{2}$ and is at a finite distance from any other point in $\mathbb{CP}^{2}$.
4 Higgs Kinetic Term
By the Higgs kinetic term we mean the combination of terms involving the Higgs field’s gauge-covariant spacetime derivatives. In our model, the Higgs field is a map $\Phi$ from spacetime to the target $\mathbb{CP}^{2}$. Its derivatives lie in the tangent space to $\mathbb{CP}^{2}$, and we use the Fubini–Study metric to construct the Higgs kinetic term, which is quadratic in these derivatives. Spacetime indices are contracted with the Minkowski metric (+,,,)(+,-,-,-)( + , - , - , - ) using the usual index raising convention.
The Higgs field in homogeneous coordinates is denoted $\Phi=(\Phi_{0},\Phi_{1},\Phi_{2})$, and in its inhomogeneous form as
$\Phi=(M,\phi_{1},\phi_{2})=(M,\phi)\,,$ (9)
where $\phi=(\phi_{1},\phi_{2})$ is a complex scalar doublet.333 All these multiplets, and the left-handed fermion doublets below, are column vectors. The Higgs field has mass dimension $1$, so the positive constant $M$ is a mass parameter. We will see later that the expansion parameter for BSM terms in our model is $\frac{v^{2}}{M^{2}}$, where $v$ is the Higgs vev, so we will need $M\gg v$.
As mentioned earlier, the $SU(3)$ symmetry is explicitly broken. This is because the point $(M,0,0)$ (equivalently $(1,0,0)$) in $\mathbb{CP}^{2}$ plays a preferred role. The isotropy group of this point is the $U(2)$ subgroup with typical element
$U=\begin{pmatrix}e^{-i\alpha/3}&0\\ 0&e^{i\alpha/6}u\end{pmatrix}$ (10)
where $u\in SU(2)$. Its action on the (inhomogeneous) Higgs field is
$\Phi\to U\Phi=(M,e^{i\alpha/2}u\phi)\,.$ (11)
In our model, as in the Standard Model, it is this $U(2)$ group that is gauged.
We introduce the $SU(2)$ and $U(1)$ gauge potentials, respectively $A_{\mu}$ and $B_{\mu}$, with $A_{\mu}$ hermitian and $B_{\mu}$ real. Using the generators $I_{a}=\frac{1}{2}\sigma_{a}$ for the Lie algebra of $SU(2)$, with $\sigma_{a}$ the Pauli matrices, we have the expansion $A_{\mu}=A_{\mu}^{a}I_{a}$ with $A_{\mu}^{a}$ real. $I_{a}$ are the weak isospin operators and the $U(1)$ generator, the hypercharge operator, is denoted $Y$. The hypercharge $y$ of a field or particle is its $Y$-eigenvalue [^4] We use the ‘half-scale’ convention for hypercharge, as in e.g. ref.[1]., and the electric charge $Q$ is the eigenvalue of $I_{3}+Y$.
In inhomogeneous coordinates, the derivatives and gauge potentials act non-trivially only on the lower doublet of fields $\phi$, i.e.
$D_{\mu}\Phi=(0,D_{\mu}\phi)$ (12)
where
$D_{\mu}\phi=\partial_{\mu}\phi-igA_{\mu}\phi-\frac{1}{2}ig^{\prime}B_{\mu}\phi\,,$ (13)
and $g$ and $g^{\prime}$ are the usual gauge coupling constants. The Higgs field $\phi$ has hypercharge $y=\frac{1}{2}$, so its complex components $\phi_{1}$ and $\phi_{2}$ have electric charges $1$ and 00, respectively.
The Higgs kinetic term is then constructed using the metric (1), with $Z$ and $dZ$ substituted by $\Phi$ and $D_{\mu}\Phi$. The result is
${\mathcal{L}}_{\rm Higgs\,Kinetic}=M^{2}\frac{(M^{2}+\bar{\phi}\phi)\overline{D_{\mu}\phi}D^{\mu}\phi-(\overline{D_{\mu}\phi}\,\phi)(\bar{\phi}D^{\mu}\phi)}{\left(M^{2}+\bar{\phi}\phi\right)^{2}}\,,$ (14)
where the prefactor $M^{2}$ is required to obtain an expression of mass dimension $4$. This kinetic term has novel features compared with the Standard Model kinetic term $\overline{D_{\mu}\phi}D^{\mu}\phi$, but reduces to it if $|\phi|$ is small compared with $M$.
5 Higgs Potential
The Higgs potential on $\mathbb{CP}^{2}$ is a real, non-negative function of the homogeneous Higgs field $\Phi$ that is unchanged by the scaling $\Phi\to\kappa\Phi$, and is $U(2)$ gauge invariant. It needs to break the $SU(3)$ symmetry of $\mathbb{CP}^{2}$, otherwise it would be constant, and it needs to have its minimum along a non-trivial, squashed 3-sphere orbit of $U(2)$, in order to spontaneously break the gauge symmetry to $U(1)_{\rm em}$ and for the Higgs mechanism to operate.
An acceptable, simple potential with these properties is
$V(\Phi)=\lambda\left(\frac{\bar{\Phi}W\Phi}{\bar{\Phi}\Phi}\right)^{2}$ (15)
with $W$ a diagonal $3\times 3$ matrix having its lower pair of entries equal, and $\lambda$ a dimensionless positive coefficient. Specifically, evaluating this for the inhomogeneous $\Phi=(M,\phi)$, and with
$W=\begin{pmatrix}\frac{1}{2}v^{2}&0&0\\ 0&-M^{2}&0\\ 0&0&-M^{2}\end{pmatrix}\,,$ (16)
we obtain
${\mathcal{L}}_{\rm Higgs\,Potential}=-V(\phi)=-\lambda M^{4}\left(\frac{\frac{1}{2}v^{2}-\bar{\phi}\phi}{M^{2}+\bar{\phi}\phi}\right)^{2}\,,$ (17)
which has mass dimension $4$. For $|\phi|\ll M$, $V(\phi)$ simplifies to its Standard Model form $\lambda\left(\frac{1}{2}v^{2}-\bar{\phi}\phi\right)^{2}$.
The minimal, vacuum value of $V$ is zero, where $\bar{\phi}\phi=\frac{1}{2}v^{2}$. This is as in the Standard Model, but here the Higgs vacuum manifold is geometrically a squashed 3-sphere, because of the form of the Higgs kinetic term (14). $V$ has a local maximum value $\frac{1}{4}\lambda v^{4}$ at the point $\phi=0$, and a further local maximum value $\lambda M^{4}$ on the 2-sphere at infinity.
As usual, the constant, vacuum field configuration is selected to be $\phi=(\phi_{1},\phi_{2})=\left(0,\frac{1}{\sqrt{2}}v\right)$, which is charge-neutral and real. The unbroken gauge group is then the usual electromagnetic $U(1)_{\rm em}$ subgroup of $U(2)$ with typical element
$u_{\rm em}=\begin{pmatrix}e^{i\beta}&0\\ 0&1\end{pmatrix}\,.$ (18)
6 Gauge Field Tensors
The gauge field tensors and their contribution to the Lagrangian are exactly as in the Standard Model. The $SU(2)$ and $U(1)$ field tensors are
$F_{\mu\nu}$ $=$ $\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}-ig[A_{\mu},A_{\nu}]\,,$ (19)
$B_{\mu\nu}$ $=$ $\partial_{\mu}B_{\nu}-\partial_{\mu}B_{\nu}\,.$ (20)
Using the expansion of the $SU(2)$ gauge potential, $A_{\mu}=A_{\mu}^{a}I_{a}$, the $SU(2)$ field tensor acquires the expansion $F_{\mu\nu}=F_{\mu\nu}^{a}I_{a}$, where
$F_{\mu\nu}^{a}=\partial_{\mu}A_{\nu}^{a}-\partial_{\nu}A_{\mu}^{a}+g\varepsilon^{abc}A_{\mu}^{b}A_{\nu}^{c}\,.$ (21)
The field tensor terms in the Lagrangian are
${\mathcal{L}}_{\rm Field\,Tensor}=-\frac{1}{4}F_{\mu\nu}^{a}F^{\mu\nu a}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}\,.$ (22)
7 Fermion Kinetic Terms
We start from the trivial $\mathbb{C}^{3}$ bundle over $\mathbb{CP}^{2}$, whose constant fibre ${\cal V}=\mathbb{C}^{3}$ is unaffected by changes in the value of the Higgs field $\Phi$, i.e. the point in $\mathbb{CP}^{2}$. We next exploit the explicit breaking of $SU(3)$ by the choice of vector $(1,0,0)\in\mathbb{C}^{3}$, which restricts the gauge group to $U(2)$ with group elements (10). This allows us to pick out the 2-dimensional subspace ${\cal V}_{L}$ of vectors in $\mathbb{C}^{3}$ whose upper component is zero, and the 1-dimensional subspace ${\cal V}_{R}$ of vectors whose lower two components are zero. Finally, using this gauge-invariant splitting, we assume that the left-handed fermion fields are (electroweak) doublets lying in ${\cal V}_{L}$ and the right-handed fermion fields are singlets lying in ${\cal V}_{R}$. $SU(2)$ acts canonically from the left on ${\cal V}_{L}$ and trivially on ${\cal V}_{R}$. The action of the $U(1)$ generator $Y$ has more freedom compatible with gauge invariance. Its eigenvalues – the hypercharges of the fermion fields – do not have to be those of the generating matrix for the group elements (10), namely $-\frac{1}{3}$ and $\frac{1}{6}$.
In the context of the Standard Model, the above remarks are a restatement of what is well known. The reason for making them is that there is an alternative possibility, exploiting the Higgs field $\Phi$. We could have split the trivial $\mathbb{C}^{3}$ bundle over $\mathbb{CP}^{2}$ into the 1-dimensional tautological bundle and the 2-dimensional tangent bundle, achieved at each point of $\mathbb{CP}^{2}$ by using the projection operators $P$ and $1-P$, as described in section 2. We could then have assumed that the left-handed fermion fields lie in the tangent bundle, and the right-handed fields in the tautological bundle. This assumption would have been compatible with retaining the full $SU(3)$ symmetry. However, it would have made the fermions dependent on the Higgs field at the kinematic level. We find this unattractive. It also makes construction of Yukawa interaction terms problematic. Instead, we have exploited the constant vector $(1,0,0)$ to produce the doublet/singlet fermion field structure. The projection operator $1-P$ will later be key to the construction of gauge-invariant Yukawa terms.
These assumptions mean that the kinetic terms (involving gauge-invariant spacetime derivatives) for the fermions are exactly as in the Standard Model, making use of the Dirac matrices $\gamma^{\mu}$ and $\gamma^{5}$. Fermion fields are doublet 4-spinors $\psi_{L}$ and singlet 4-spinors $\psi_{R}$ satisfying, respectively, $\gamma^{5}\psi_{L}=-\psi_{L}$ and $\gamma^{5}\psi_{R}=\psi_{R}$.
For the lightest, first-generation quark pair, the u-quark and d-quark, the fields are the left-handed doublet $q_{L}=(u_{L},d_{L})$ and the two right-handed singlets $u_{R}$ and $d_{R}$. The second- and third-generation quark pairs are similar and we do not write their Lagrangians explicitly. The quark kinetic terms are
${\mathcal{L}}_{\rm Quark\,Kinetic}=\overline{q_{L}}\,i\gamma^{\mu}D_{\mu}q_{L}+\overline{u_{R}}\,i\gamma^{\mu}D_{\mu}u_{R}+\overline{d_{R}}\,i\gamma^{\mu}D_{\mu}d_{R}\,,$ (23)
where the different gauge-covariant derivatives are
$D_{\mu}q_{L}$ $=$ $\partial_{\mu}q_{L}-igA_{\mu}q_{L}-ig^{\prime}y_{L}^{q}B_{\mu}q_{L}\,,$ (24)
$D_{\mu}u_{R}$ $=$ $\partial_{\mu}u_{R}-ig^{\prime}y_{R}^{u}B_{\mu}u_{R}\,,$ (25)
$D_{\mu}d_{R}$ $=$ $\partial_{\mu}d_{R}-ig^{\prime}y_{R}^{d}B_{\mu}d_{R}$ (26)
and the quark hypercharges are $(y_{L}^{q},y_{R}^{u},y_{R}^{d})=(\frac{1}{6},\frac{2}{3},-\frac{1}{3})$. Note that the overbar on the fermion fields in (23) is the Dirac conjugate, including $\gamma^{0}$.
For the first-generation lepton pair, the electron and its neutrino, the fields are the doublet $l_{L}=(\nu_{L},e_{L})$ and the one singlet $e_{R}$. For the second- and third-generation lepton pairs we again do not write the Lagrangians explicitly. The lepton kinetic terms are
${\mathcal{L}}_{\rm Lepton\,Kinetic}=\overline{l_{L}}\,i\gamma^{\mu}D_{\mu}l_{L}+\overline{e_{R}}\,i\gamma^{\mu}D_{\mu}e_{R}\,,$ (27)
with the gauge-covariant derivatives
$D_{\mu}l_{L}$ $=$ $\partial_{\mu}l_{L}-igA_{\mu}l_{L}-ig^{\prime}y_{L}^{l}B_{\mu}l_{L}\,,$ (28)
$D_{\mu}e_{R}$ $=$ $\partial_{\mu}e_{R}-ig^{\prime}y_{R}^{e}B_{\mu}e_{R}$ (29)
and the lepton hypercharges $(y_{L}^{l},y_{R}^{e})=(-\frac{1}{2},-1)$.
These fermions have electric charges $(Q^{u},Q^{d},Q^{\nu},Q^{e})=(\frac{2}{3}.-\frac{1}{3},0,-1)$.
8 Yukawa Terms
The Yukawa terms in the Standard Model generate gauge-invariant masses for the fermions. They couple the Higgs field to fermion bilinears, and the masses generated are proportional to a Yukawa coupling times the Higgs vev $v$. We first recall the two ways for contracting the left-handed fermion doublets with the Higgs doublet field $\phi$, in order to have $SU(2)$ singlets. Let us illustrate this for the left-handed quark doublet $q_{L}$. In components,
$q_{L}=(u_{L},d_{L})\quad{\rm and}\quad\phi=(\phi_{1},\phi_{2})\,.$ (30)
The first contraction is the obvious $\overline{q_{L}}\,\phi$, and the Yukawa term obtained from this is
$-\lambda_{d}\,\overline{q_{L}}\,\phi\,d_{R}=-\lambda_{d}(\overline{u_{L}}\,\phi_{1}\,d_{R}+\overline{d_{L}}\,\phi_{2}\,d_{R})\,,$ (31)
which is a $U(2)$ singlet because of the quark and Higgs field hypercharge assignments. Only a d-quark mass is generated through this term in the Higgs vacuum, $(\phi_{1},\phi_{2})=\left(0,\frac{1}{\sqrt{2}}v\right)$. To generate a u-quark mass, one needs to exploit the conjugate of the Higgs doublet, $\phi_{C}=i\sigma_{2}\phi^{*}=(\phi_{2}^{*},-\phi_{1}^{*})$. This is also a doublet of $SU(2)$, like $\phi$, but has the opposite hypercharge. The second contraction with $q_{L}$ is then $\overline{q_{L}}\,\phi_{C}$, and from this one obtains the Yukawa term
$-\lambda_{u}\,\overline{q_{L}}\,\phi_{C}\,u_{R}=-\lambda_{u}(\overline{u_{L}}\,\phi_{2}^{*}\,u_{R}-\overline{d_{L}}\,\phi_{1}^{*}\,u_{R})\,.$ (32)
$\lambda_{d}$ and $\lambda_{u}$ are the Yukawa coupling constants that combine with the Higgs vev to give the d- and u-quark masses $\frac{1}{\sqrt{2}}\lambda_{d}v$ and $\frac{1}{\sqrt{2}}\lambda_{u}v$.
For the leptons there is just one Yukawa term, giving the electron a mass. Since $e_{L}$ is the lower component of its doublet, the required term is
$-\lambda_{e}\,\overline{l_{L}}\,\phi\,e_{R}=-\lambda_{e}(\overline{\nu_{L}}\,\phi_{1}\,e_{R}+\overline{e_{L}}\,\phi_{2}\,e_{R})\,,$ (33)
the analogue of the term giving the d-quark mass. The electron mass is $\frac{1}{\sqrt{2}}\lambda_{e}v$. The neutrino is left-handed and massless, with no right-handed field, so there is no need to use $\phi_{C}$ again.
It is rather mysterious that the Higgs field occurs in two different ways, but it is straightforward to verify directly that $\phi$ and $\phi_{C}$ transform the same way under the left-action of $SU(2)$. This can also be understood by considering left multiplication by the $SU(2)$ matrix
$u=\begin{pmatrix}a&b\\ -b^{*}&a^{*}\end{pmatrix}$ (34)
(with $|a|^{2}+|b|^{2}=1$) on the $2\times 2$ Higgs field matrix
$\phi_{m}=\begin{pmatrix}\phi_{2}^{*}&\phi_{1}\\ -\phi_{1}^{*}&\phi_{2}\end{pmatrix}\,,$ (35)
with columns $\phi_{C}$ and $\phi$. $\phi_{m}$ is an $SU(2)$ matrix times the scalar multiple $\sqrt{|\phi_{1}|^{2}+|\phi_{2}|^{2}}$. The product $\phi^{\prime}_{m}=u\phi_{m}$ is consequently another $SU(2)$ matrix times the same multiple and has the same form, with $(\phi_{1},\phi_{2})$ replaced by
$(\phi^{\prime}_{1},\phi^{\prime}_{2})=(a\phi_{1}+b\phi_{2},-b^{*}\phi_{1}+a^{*}\phi_{2})\,.$ (36)
Matrix multiplication by $u$ acts the same way on the left and right columns of $\phi_{m}$, and therefore $\phi$ and $\phi_{C}$ transform the same way under $SU(2)$.
We now turn to our $\mathbb{CP}^{2}$ model. There is still no unification of $\phi$ and $\phi_{C}$, and two distinct types of Yukawa term are needed. We will use the homogeneous Higgs field $\Phi$ in our Yukawa terms, and must ensure that the result is gauge invariant and unchanged by rescaling by $\kappa$. The useful scale-invariant objects derived from $\Phi$ are the projectors
$P=\frac{\Phi\bar{\Phi}}{\bar{\Phi}\Phi}\quad{\rm and}\quad 1-P=\frac{\bar{\Phi}\Phi-\Phi\bar{\Phi}}{\bar{\Phi}\Phi}\,.$ (37)
In terms of the components of the inhomogeneous Higgs field $\Phi=(M,\phi_{1},\phi_{2})$, these projectors are the $3\times 3$ matrices
$P=\frac{1}{M^{2}+|\phi_{1}|^{2}+|\phi_{2}|^{2}}\begin{pmatrix}M^{2}&M\phi_{1}^{*}&M\phi_{2}^{*}\\ M\phi_{1}&|\phi_{1}|^{2}&\phi_{1}\phi_{2}^{*}\\ M\phi_{2}&\phi_{2}\phi_{1}^{*}&|\phi_{2}|^{2}\end{pmatrix}$ (38)
and
$1-P=\frac{1}{M^{2}+|\phi_{1}|^{2}+|\phi_{2}|^{2}}\begin{pmatrix}|\phi_{1}|^{2}+|\phi_{2}|^{2}&-M\phi_{1}^{*}&-M\phi_{2}^{*}\\ -M\phi_{1}&M^{2}+|\phi_{2}|^{2}&-\phi_{1}\phi_{2}^{*}\\ -M\phi_{2}&-\phi_{2}\phi_{1}^{*}&M^{2}+|\phi_{1}|^{2}\end{pmatrix}.$ (39)
Now recall that in our model the left-handed quark doublet and the right-handed d-quark singlet are (column) 3-vector fields $q_{L}=(0,u_{L},d_{L})$ and $(d_{R},0,0)$. Acting with $1-P$ on the d-quark singlet gives the 3-vector
$\frac{1}{M^{2}+|\phi_{1}|^{2}+|\phi_{2}|^{2}}\biggl{(}(|\phi_{1}|^{2}+|\phi_{2}|^{2})d_{R},\,-M\phi_{1}\,d_{R},\,-M\phi_{2}\,d_{R}\biggr{)}.$ (40)
Contracting with $q_{L}$, and including an appropriate coupling constant, gives the Yukawa term
$\frac{-\Lambda_{d}M^{2}}{M^{2}+|\phi_{1}|^{2}+|\phi_{2}|^{2}}(\overline{u_{L}}\,\phi_{1}\,d_{R}+\overline{d_{L}}\,\phi_{2}\,d_{R})\,,$ (41)
where we have included an extra factor of $M$ so that this term has mass dimension 4. The Yukawa coupling $\Lambda_{d}$ is dimensionless and close to its value in the Standard Model. When the Higgs field acquires its vev, the d-quark mass becomes
$m_{d}=\frac{1}{\sqrt{2}}\frac{M^{2}}{M^{2}+\frac{1}{2}v^{2}}\Lambda_{d}v\,.$ (42)
The Yukawa term for the lepton doublet is similar, giving an electron mass
$m_{e}=\frac{1}{\sqrt{2}}\frac{M^{2}}{M^{2}+\frac{1}{2}v^{2}}\Lambda_{e}v\,.$ (43)
Note that the projector $1-P$ acting on a right-handed fermion singlet produces a doublet and a singlet. The doublet plays the role in our $\mathbb{CP}^{2}$ model that the product of $\phi$ with the right-handed singlet does in the Standard Model, and its contraction with a left-handed fermion doublet is $U(2)$ invariant. One could generate further Yukawa terms by using the projected singlet, but these would be of higher order in $\phi$.
For the second type of Yukawa term, giving the u-quark its mass, we need to replace $\phi$ by its conjugate $\phi_{C}$, and use this to construct the conjugate projector $1-P_{C}$. The right-handed u-quark singlet is the 3-vector $(u_{R},0,0)$. Acting with $1-P_{C}$ on this, and contracting with $q_{L}$, gives the Yukawa term
$\frac{-\Lambda_{u}M^{2}}{M^{2}+|\phi_{1}|^{2}+|\phi_{2}|^{2}}(\overline{u_{L}}\,\phi_{2}^{*}\,u_{R}-\overline{d_{L}}\,\phi_{1}^{*}\,u_{R})\,.$ (44)
The u-quark mass is then
$m_{u}=\frac{1}{\sqrt{2}}\frac{M^{2}}{M^{2}+\frac{1}{2}v^{2}}\Lambda_{u}v\,.$ (45)
The total contribution of the Yukawa terms to the Lagrangian of the $\mathbb{CP}^{2}$ model is
${\mathcal{L}}_{\rm Yukawa}$ $=$ $-\,\frac{M^{2}}{M^{2}+|\phi_{1}|^{2}+|\phi_{2}|^{2}}\Bigl{[}\Lambda_{d}(\overline{u_{L}}\,\phi_{1}\,d_{R}+\overline{d_{L}}\,\phi_{2}\,d_{R})$ (46)
$\quad+\Lambda_{u}(\overline{u_{L}}\,\phi_{2}^{*}\,u_{R}-\overline{d_{L}}\,\phi_{1}^{*}\,u_{R})+\Lambda_{e}(\overline{\nu_{L}}\,\phi_{1}\,e_{R}+\overline{e_{L}}\,\phi_{2}\,e_{R})\Bigr{]}.$
9 Phenomenology Beyond the Standard Model
The Lagrangian of our proposed $\mathbb{CP}^{2}$ SMEFT is
${\mathcal{L}}$ $=$ ${\mathcal{L}}_{\rm Field\,Tensor}+{\mathcal{L}}_{\rm Higgs\,Kinetic}+{\mathcal{L}}_{\rm Higgs\,Potential}$ (47)
$\qquad+\,{\mathcal{L}}_{\rm Quark\,Kinetic}+{\mathcal{L}}_{\rm Lepton\,Kinetic}+{\mathcal{L}}_{\rm Yukawa}\,,$
with its contributions defined in sections 5 – 8 above. Some of the apparent differences from the Standard Model can be removed by a recalibration of parameters and rescaling of the fields. But there remain some significant modifications to SM relations, and we will use these together with recent experimental results as reviewed by the Particle Data Group (PDG) [7] to place constraints on the BSM mass parameter $M$. Also, when the denominator factors are expanded out, there arise several dimension-6 terms that are Standard Model terms multiplied by $\bar{\phi}\phi$. In the following, we will explore the consequences of the $\mathbb{CP}^{2}$ SMEFT for (i) gauge boson masses, (ii) Higgs cubic and quartic couplings, (iii) Yukawa couplings. Constraints on $M$ currently come mainly from the $W$-boson to $Z$-boson mass ratio.
9.1 The $W$-boson to $Z$-boson mass ratio
The gauge boson masses are determined using the terms in ${\mathcal{L}}$ involving the gauge field tensor and covariant derivatives of the Higgs field. Most interesting is the novel contribution $(\overline{D_{\mu}\phi}\,\phi)(\bar{\phi}D^{\mu}\phi)$ to the Higgs kinetic term.
The gauge couplings $g$ and $g^{\prime}$ are the same as in the Standard Model, as they control the tree-level couplings of gauge bosons to fermions. The weak mixing angle $\theta_{w}$ is therefore also the same; it is defined, as usual, via
$\cos\theta_{w}=\frac{g}{\sqrt{g^{2}+g^{\prime 2}}}\,,\quad\sin\theta_{w}=\frac{g^{\prime}}{\sqrt{g^{2}+g^{\prime 2}}}\,.$ (48)
The field tensor term in the Lagrangian (47) has its canonical form, so the (squared) spacetime derivatives of the gauge boson fields have canonical normalisation. The gauge boson masses are then extracted from the gauge-covariant Higgs kinetic term (14) by setting the Higgs field equal to its vacuum value $\phi=\left(0,\frac{1}{\sqrt{2}}v\right)$, and ignoring the derivatives of $\phi$. The two contributions (without their prefactors) are
$\overline{D_{\mu}\phi}D^{\mu}\phi$ $=$ $\frac{v^{2}}{8}\left[g^{2}(A_{\mu}^{1})^{2}+g^{2}(A_{\mu}^{2})^{2}+(gA_{\mu}^{3}-g^{\prime}B_{\mu})^{2}\right]\,,$
$(\overline{D_{\mu}\phi}\,\phi)(\bar{\phi}D^{\mu}\phi)$ $=$ $\frac{v^{4}}{16}(gA_{\mu}^{3}-g^{\prime}B_{\mu})^{2}\,,$ (49)
(where $(A_{\mu}^{1})^{2}$ is shorthand for $A_{\mu}^{1}A^{1\mu}$, etc.). The second of these does not appear at all in the Standard Model Lagrangian. It is the term denoted ${\cal Q}_{\phi D}$ [8] or ${\cal O}_{\phi D}$ [9] in the classification of dimension-6 SMEFT operators, and denoted ${\cal O}_{HD}$ [10, 11] in the analysis of constraints on their coefficients.
As usual, we define the $Z$-boson field and photon field to be
$Z_{\mu}=\frac{1}{\sqrt{g^{2}+g^{\prime 2}}}(gA_{\mu}^{3}-g^{\prime}B_{\mu})\,,\quad A_{\mu}=\frac{1}{\sqrt{g^{2}+g^{\prime 2}}}(g^{\prime}A_{\mu}^{3}+gB_{\mu})\,,$ (50)
or equivalently
$Z_{\mu}=\cos\theta_{w}\,A_{\mu}^{3}-\sin\theta_{w}\,B_{\mu}\,,\quad A_{\mu}=\sin\theta_{w}\,A_{\mu}^{3}+\cos\theta_{w}\,B_{\mu}\,.$ (51)
The $W$-boson fields are $W_{\mu}^{\pm}=\frac{1}{\sqrt{2}}(A_{\mu}^{1}\mp A_{\mu}^{2})$. It is not surprising that both gauge-boson mass terms (49) have no contribution from the photon field. The photon remains massless because of the unbroken $U(1)_{\rm em}$ symmetry.
The combined gauge-boson mass term, including the prefactors, is
$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\biggl{[}\left(M^{2}+\frac{1}{2}v^{2}\right)\frac{v^{2}}{8}\left[g^{2}(A_{\mu}^{1})^{2}+g^{2}(A_{\mu}^{2})^{2}+(g^{2}+g^{\prime 2})Z_{\mu}^{2}\right]$
$\qquad\qquad\quad\qquad\qquad\qquad-\frac{v^{4}}{16}(g^{2}+g^{\prime 2})Z_{\mu}^{2}\biggr{]}.$ (52)
$v^{4}Z_{\mu}^{2}$ cancels here, and what remains can be reexpressed as
$\frac{1}{2}\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\left(\frac{gv}{2}\right)^{2}\left[\left(M^{2}+\frac{1}{2}v^{2}\right)((A_{\mu}^{1})^{2}+(A_{\mu}^{2})^{2})+\frac{M^{2}}{\cos^{2}\theta_{w}}Z_{\mu}^{2}\right].$ (53)
From this we read off the gauge boson masses
$m_{W}=\frac{M}{\sqrt{M^{2}+\frac{1}{2}v^{2}}}\left(\frac{gv}{2}\right)\,,\quad m_{Z}=\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)}\left(\frac{gv}{2}\right)\frac{1}{\cos\theta_{w}}\,.$ (54)
From here on, we work to leading order in $\frac{v^{2}}{M^{2}}$, as our estimates below show that this quantity is less than $0.01$. The masses become
$m_{W}=\left(1-\frac{v^{2}}{4M^{2}}\right)\left(\frac{gv}{2}\right)\,,\quad m_{Z}=\left(1-\frac{v^{2}}{2M^{2}}\right)\left(\frac{gv}{2}\right)\frac{1}{\cos\theta_{w}}\,.$ (55)
The Higgs vev $v$ can be expressed in terms of the conventional $v_{\rm SM}=246.22$ GeV to match the $Z$-boson mass formula in the Standard Model. To do this, we define
$v_{\rm SM}=\left(1-\frac{v^{2}}{2M^{2}}\right)v$ (56)
and find
$m_{W}=\left(1+\frac{v_{\rm SM}^{2}}{4M^{2}}\right)\left(\frac{gv_{\rm SM}}{2}\right)\,,\quad m_{Z}=\left(\frac{gv_{\rm SM}}{2}\right)\frac{1}{\cos\theta_{w}}\,.$ (57)
The key result here is that the squared $W$-boson to $Z$-boson mass relation is
$m_{W}^{2}=\left(1+\frac{v_{\rm SM}^{2}}{2M^{2}}\right)m_{Z}^{2}\cos^{2}\theta_{w}\,.$ (58)
This is a small modification of the Standard Model tree level relation $m_{W}^{2}=m_{Z}^{2}\cos^{2}\theta_{w}$ in the direction of increasing the $W$-boson mass towards the $Z$-boson mass.
There are SM loop corrections to this tree level relation, but we can assume that these are essentially the same in the $\mathbb{CP}^{2}$ SMEFT. The $\mathbb{CP}^{2}$ SMEFT therefore predicts a BSM “new physics” factor
$\rho_{0}=1+\frac{v_{\rm SM}^{2}}{2M^{2}}\,.$ (59)
Current data [7] gives $\rho_{0}=1.00031\pm 0.00019$ if the combined CERN and Fermilab D0 result of 80.360 GeV for the $W$-boson mass is used, i.e. slightly higher than the SM prediction of $\rho_{0}=1$ but of little statistical significance. Any BSM correction to the SM is therefore small, but appears to be in the direction of an increase of the $W$-boson mass relative to the $Z$-boson mass, as we find in the $\mathbb{CP}^{2}$ SMEFT. At a confidence level of $2\sigma$, $\rho_{0}<1.00031+0.00038=1.00069$ and formula (59) implies that with the same confidence, the $\mathbb{CP}^{2}$ SMEFT requires $M>27\,v_{\rm SM}\simeq 6.6$ TeV.
The Fermilab CDF result for the $W$-boson mass, based on a recent reanalysis of older data [12], is 80.434 GeV, with an error estimate making this value $5\sigma$ higher than the combined CERN/D0 result and $7\sigma$ higher than the SM prediction. If we use this, then $\rho_{0}\simeq 1.002$, and from (59) we deduce that $M\simeq 4$ TeV. Within our model, this estimate for $M$ is not ruled out by the Yukawa couplings, to be discussed below.
9.2 Higgs-boson mass and couplings
Here we consider the terms of the Lagrangian involving the Higgs field alone, leaving out couplings of the Higgs field to the gauge potentials and fermions.
In the SM, these terms are
${\mathcal{L}}_{\rm Higgs}=\overline{\partial_{\mu}\phi}\partial^{\mu}\phi-\lambda\left(\frac{1}{2}v_{\rm SM}^{2}-\overline{\phi}\phi\right)^{2}\,,$ (60)
and in unitary gauge
$\phi=\frac{1}{\sqrt{2}}(0,v_{\rm SM}+H)\,,$ (61)
where $v_{\rm SM}$ is the SM Higgs vev and $H$ the real Higgs boson field. Expanding out, we obtain
${\mathcal{L}}_{\rm Higgs}=\frac{1}{2}\partial_{\mu}H\partial^{\mu}H-\lambda v_{\rm SM}^{2}H^{2}-\lambda v_{\rm SM}H^{3}-\frac{1}{4}\lambda H^{4}\,.$ (62)
A standard notation for this is
${\mathcal{L}}_{\rm Higgs}=\frac{1}{2}\partial_{\mu}H\partial^{\mu}H-\frac{1}{2}m_{H}^{2}H^{2}-\frac{1}{6}\lambda_{3}H^{3}-\frac{1}{24}\lambda_{4}H^{4}\,,$ (63)
so $m_{H}^{2}=2\lambda v_{\rm SM}^{2}$ is the Higgs boson mass, $\lambda_{3}=6\lambda v_{\rm SM}=3\frac{m_{H}^{2}}{v_{\rm SM}}$ is the cubic Higgs coupling, and $\lambda_{4}=6\lambda=3\frac{m_{H}^{2}}{v_{\rm SM}^{2}}$ is the quartic Higgs coupling. As $m_{H}$ and $v_{\rm SM}$ are precisely known, the cubic and quartic Higgs couplings have precisely predicted SM values.
In the $\mathbb{CP}^{2}$ SMEFT, terms in ${\mathcal{L}}_{\rm Higgs}$ acquire an additional denominator factor $(1+\frac{\overline{\phi}\phi}{M^{2}})^{2}$. Working to leading order, i.e. up to dimension 6 terms, we can replace this by a numerator factor $(1-2\frac{\overline{\phi}\phi}{M^{2}})$. Then, going to unitary gauge and writing
$\phi=\frac{1}{\sqrt{2}}(0,v+h)\,,$ (64)
we find from the kinetic and potential contributions (14) and (17) that
${\mathcal{L}}_{\rm Higgs}$ $=$ $\frac{1}{2}\left(1-\frac{v^{2}}{M^{2}}-\frac{2v}{M^{2}}h-\frac{1}{M^{2}}h^{2}\right)\partial_{\mu}h\partial^{\mu}h$ (65)
$\ -\frac{1}{4}\lambda\left(1-\frac{v^{2}}{M^{2}}-\frac{2v}{M^{2}}h-\frac{1}{M^{2}}h^{2}\right)(4v^{2}h^{2}+4vh^{3}+h^{4}).$
Note that this includes terms where $\frac{1}{2}\partial_{\mu}h\partial^{\mu}h$ is multiplied by $h$ and $h^{2}$. These are small, momentum dependent corrections to the cubic and quartic Higgs couplings. From now on, we drop these, leaving, after collecting terms,
${\mathcal{L}}_{\rm Higgs}$ $=$ $\frac{1}{2}\left(1-\frac{v^{2}}{M^{2}}\right)\partial_{\mu}h\partial^{\mu}h-\lambda v^{2}\left(1-\frac{v^{2}}{M^{2}}\right)h^{2}$ (66)
$\quad-\lambda v\left(1-\frac{3v^{2}}{M^{2}}\right)h^{3}-\frac{1}{4}\lambda\left(1-\frac{13v^{2}}{M^{2}}\right)h^{4}+O(h^{5})\,.$
We need to rescale $h$ to reproduce the canonical kinetic term for the Higgs boson field. So we define
$H=\left(1-\frac{v^{2}}{2M^{2}}\right)h$ (67)
and obtain (to leading order in $\frac{v^{2}}{M^{2}}$)
${\mathcal{L}}_{\rm Higgs}$ $=$ $\frac{1}{2}\partial_{\mu}H\partial^{\mu}H-\lambda v^{2}H^{2}$ (68)
$\quad-\lambda v\left(1-\frac{3v^{2}}{2M^{2}}\right)H^{3}-\frac{1}{4}\lambda\left(1-\frac{11v^{2}}{M^{2}}\right)H^{4}\,.$
$m_{H}^{2}=2\lambda v^{2}$ is therefore the Higgs boson mass, and we can eliminate $\lambda$ in favour of $m_{H}$ in the cubic and quartic terms to obtain
${\mathcal{L}}_{\rm Higgs}$ $=$ $\frac{1}{2}\partial_{\mu}H\partial^{\mu}H-\lambda v^{2}H^{2}$ (69)
$\quad-\frac{m_{H}^{2}}{2v}\left(1-\frac{3v^{2}}{2M^{2}}\right)H^{3}-\frac{m_{H}^{2}}{8v^{2}}\left(1-\frac{11v^{2}}{M^{2}}\right)H^{4}\,.$
Writing this again in the form (63), we read off that
$\lambda_{3}=\frac{3m_{H}^{2}}{v}\left(1-\frac{3v^{2}}{2M^{2}}\right)\quad{\rm and}\quad\lambda_{4}=\frac{3m_{H}^{2}}{v^{2}}\left(1-\frac{11v^{2}}{M^{2}}\right)\,.$ (70)
To compare with the SM couplings we need to make the transformation (56) from $v$ to $v_{\rm SM}$. To leading order,
$\lambda_{3}=\frac{3m_{H}^{2}}{v_{\rm SM}}\left(1-\frac{2v_{\rm SM}^{2}}{M^{2}}\right)\quad{\rm and}\quad\lambda_{4}=\frac{3m_{H}^{2}}{v_{\rm SM}^{2}}\left(1-\frac{12v_{\rm SM}^{2}}{M^{2}}\right)\,.$ (71)
Then, writing these couplings as [13]
$\lambda_{3}=\frac{3m_{H}^{2}}{v_{\rm SM}}(1+\kappa_{3})\quad{\rm and}\quad\lambda_{4}=\frac{3m_{H}^{2}}{v_{\rm SM}^{2}}(1+\kappa_{4})$ (72)
we read off the BSM modifications
$\kappa_{3}=-\frac{2v_{\rm SM}^{2}}{M^{2}}\quad{\rm and}\quad\kappa_{4}=-\frac{12v_{\rm SM}^{2}}{M^{2}}\,.$ (73)
The $\mathbb{CP}^{2}$ SMEFT therefore predicts that both these parameters are negative and in the ratio $\kappa_{4}=6\kappa_{3}$.
Unfortunately, the cubic and quartic Higgs couplings are not yet determined experimentally. There are only weak upper bounds on their strengths. In future, using the HL-LHC at CERN, it is hoped that these couplings can be measured and compared with SM predictions. Even then, the small modifications $\kappa_{3}$ and $\kappa_{4}$ to the SM predictions will probably remain out of reach. So these couplings give no useful constraint on $M$. However, the discussion here has clarified the need for the rescaling of the Higgs boson field $h$ as well as the Higgs vev $v$. These rescalings have an important effect on the Yukawa couplings, which we calculate next.
9.3 Higgs-fermion couplings
The $\mathbb{CP}^{2}$ SMEFT modifies the SM relationship between fermion masses and the coupling of fermions to the Higgs boson. As we will see, this modification is by a universal factor for all fermion types. The coupling controls both the production and decay strengths for the Higgs boson, and has been accurately measured for the three fermions of the third generation – the t-quark and b-quark, and the $\tau$-lepton. Currently, the SM relationship is confirmed, within experimental uncertainties, so this places limits on BSM modifications, and hence a lower bound on $M$ in the $\mathbb{CP}^{2}$ SMEFT.
The kinetic terms for fermions have their canonical form in ${\mathcal{L}}$, so we need only investigate the Yukawa terms. To determine the fermion masses and coupling we need to retain only the terms up to linear order in the Higgs field $H$. As before, we convert the denominator factor in ${\mathcal{L}}_{\rm Yukawa}$ to a numerator factor and obtain for any fermion $f$,
${\mathcal{L}}_{\rm Yukawa}$ $=$ $-\frac{1}{\sqrt{2}}\Lambda_{f}\overline{f}f\left(1-\frac{v^{2}}{2M^{2}}-\frac{v}{M^{2}}h-\frac{1}{2M^{2}}h^{2}\right)(v+h)$ (74)
$=$ $-\frac{1}{\sqrt{2}}\Lambda_{f}\overline{f}f\left(\left(1-\frac{v^{2}}{2M^{2}}\right)v+\left(1-\frac{3v^{2}}{2M^{2}}\right)h+O(h^{2})\right).$
We next rescale $v$ and $h$ by the common factor $\left(1-\frac{v^{2}}{2M^{2}}\right)$ to obtain the normalised quantities $v_{\rm SM}$ and $H$. To linear order in $H$, the Yukawa term becomes
${\mathcal{L}}_{\rm Yukawa}=-\frac{1}{\sqrt{2}}\Lambda_{f}\overline{f}f\left(v_{\rm SM}+\left(1-\frac{v_{\rm SM}^{2}}{M^{2}}\right)H\right)\,.$ (75)
The mass of the fermion $f$ and its Yukawa coupling to $H$ are
$m_{f}=\frac{1}{\sqrt{2}}\Lambda_{f}v_{\rm SM}\quad{\rm and}\quad y_{f}=\frac{1}{\sqrt{2}}\Lambda_{f}\left(1-\frac{v_{\rm SM}^{2}}{M^{2}}\right)\,,$ (76)
giving the universal BSM ratio
$\kappa_{f}\equiv\frac{y_{f}}{m_{f}}v_{\rm SM}=1-\frac{v_{\rm SM}^{2}}{M^{2}}\,,$ (77)
a slight decrease compared to the SM value $1$.
The measured values of this ratio [7] are $\kappa_{t}=1.0\pm 0.1$, $\kappa_{b}=1.0\pm 0.15$, $\kappa_{\tau}=0.9\pm 0.1$, which are all consistent with the SM. For the sake of argument, suppose that more precise measurements give a universal value $\kappa_{f}=0.95$. Then $M=$ O(1 TeV). More realistically, the current data imply that $M$ is O(1 TeV) or larger in the $\mathbb{CP}^{2}$ SMEFT. This lower bound is weaker than that obtained from the $W$-boson to $Z$-boson mass ratio, so it is the mass ratio that currently is most restrictive. Note that this lower bound does not conflict with the CDF result for the $W$-boson to $Z$-boson mass ratio.
10 Conclusions
We have proposed a highly constrained SMEFT, where the Higgs field takes its value in the complex projective plane $\mathbb{CP}^{2}$ rather than in $\mathbb{C}^{2}$. This is a truly nonlinear modification of the Higgs sector of the Standard Model. The Lagrangian uses the Fubini–Study metric on $\mathbb{CP}^{2}$, and makes some use of its $SU(3)$ symmetry. The only new parameter is a BSM mass scale $M$, and the modification of the Standard Model is small provided $M\gg v$, with $v$ the Higgs vev. The proposed model is fully consistent with $U(2)$ gauge invariance; however, the custodial $SO(4)$ global symmetry of the Higgs sector of the Standard Model is mildly broken, because the $U(2)$ orbits on $\mathbb{CP}^{2}$ are squashed rather than round 3-spheres. The key prediction is that, as a consequence of this squashed geometry, the $W$-boson mass is closer to the $Z$-boson mass than in the Standard Model, while the weak mixing angle $\theta_{w}$ remains unchanged. Yukawa couplings of the Higgs boson to fermions are also modified by a universal factor. Almost all current data are consistent with the Standard Model, so, as for other BSM scenarios, there is no compelling evidence yet in favour of the $\mathbb{CP}^{2}$ SMEFT. The data place a lower bound of O(7 TeV) on $M$, unless the CDF measurement of the $W$-boson mass turns out to be correct, in which case some physics beyond the Standard Model would be needed to explain this, and within the $\mathbb{CP}^{2}$ SMEFT, $M$ would be about 4 TeV. More precise and consistent data for the gauge boson masses, Yukawa couplings and the Higgs boson self-couplings would all be helpful.
The $\mathbb{CP}^{2}$ SMEFT has no obvious UV-completion. Possibly, it can be extended to a UV-complete theory where the $SU(3)$ symmetry is dynamically realised and spontaneously broken to $U(1)$ in two steps, but that may require several additional gauge and Higgs fields, and additional fermions too. The simplest idea would be to first break the $SU(3)$ symmetry with a complex triplet Higgs field, but the unbroken gauge group would then be $SU(2)$ rather than $U(2)$, and the projective space $\mathbb{CP}_{2}$ would not appear. There are several proposed unified models with gauge group $SU(3)$ (or products of $SU(3)$ with another group) acting linearly on multiple Higgs fields, e.g. [14, 15, 16, 17, 18, 19]. These models mostly have substantially more gauge bosons, scalar particles and fermions than the Standard Model, some with exotic charges. These models do not seem to be equivalent to, or extensions of, the $\mathbb{CP}^{2}$ SMEFT proposed here, although the gauge boson mass formulae in [18] are very similar to our formulae (55).
Acknowledgements
I am grateful to Mia West and Eetu Loisa for stimulating discussions about SMEFTs, initiated at the STFC HEP Forum: Completing the Higgs-saw puzzle, Cosener’s House, Abingdon, November ’23. I also thank Alex Mitov for comments. Additionally, I wish to thank the anonymous referee for several constructive suggestions which have been incorporated in this revised paper. This work has been partially supported by STFC consolidated grant ST/P000681/1.

### v3

@@
 The Lagrangian has terms involving: (i) Gauge-covariant derivatives of the Higgs field, (ii) a Higgs potential, (iii) Gauge field tensors, (iv) Dirac (kinetic) terms for fermions, (v) Yukawa terms coupling the Higgs field and fermions.
-We will discuss these in turn, and then explore some phenomenological consequences, in particular, for the $W$-boson to $Z$-boson mass ratio, and for the Yukawa couplings of the Higgs boson.
+We will discuss these in turn, and then explore some phenomenological consequences, in particular, for the $W$-boson to $Z$-boson mass ratio, for the couplings of the $W$- and $Z$-bosons to the physical Higgs boson [^2] The earlier published version, JHEP 04 (2025) 180, did not include this topic., and for the Yukawa couplings of the Higgs boson.
 These will lead to constraints on the BSM mass parameter $M$.
@@
 Geometrically, $\mathbb{CP}^{2}$ is the manifold of complex lines [^3] We use the mathematician’s notion of complex line $\mathbb{C}$ to denote what is more commonly called the complex plane.
-All lines considered here are lines through the origin of $\mathbb{C}^{3}$.
-in $\mathbb{C}^{3}$.
+through the origin in $\mathbb{C}^{3}$.
 We denote a non-zero point in $\mathbb{C}^{3}$ by the (column) vector $Z=(Z_{0},Z_{1},Z_{2})$ and the scaling action is $Z\to\kappa Z$ with $\kappa\in\mathbb{C}-\{0\}$.
@@
 The generic orbits of $U(2)$ on $\mathbb{CP}^{2}$ are 3-spheres, as will be clarified in the next section.
-Later, we will be using the projections of a generic vector $\Psi\in\mathbb{C}^{3}$, defined at location $Z\in\mathbb{C}^{3}$, on to the line through $Z$ and on to its complement, which is a copy of the tangent space to $\mathbb{CP}^{2}$.
+Later, we will be using the projections of a generic vector $\Psi\in\mathbb{C}^{3}$, defined at location $Z\in\mathbb{C}^{3}$, on to the line through the origin and $Z$ and on to its complement, which is a copy of the tangent space to $\mathbb{CP}^{2}$.
 These projections are respectively
@@
 Because these projections are unaffected by rescaling $Z$ by $\kappa$, they are defined at location $Z$ on $\mathbb{CP}^{2}$.
-More precisely, over $\mathbb{CP}^{2}$ there is its 1-dimensional tautological bundle, whose fibre at a point $Z\in\mathbb{CP}^{2}$ is the line in $\mathbb{C}^{3}$ through $Z$, and its 2-dimensional tangent bundle.
+More precisely, over $\mathbb{CP}^{2}$ there is its 1-dimensional tautological bundle, whose fibre at a point $Z\in\mathbb{CP}^{2}$ is the line in $\mathbb{C}^{3}$ through the origin and $Z$, and its 2-dimensional tangent bundle.
 The 3-dimensional, trivial $\mathbb{C}^{3}$ bundle over $\mathbb{CP}^{2}$ is the direct sum of these, so an element $\Psi$ in the trivial $\mathbb{C}^{3}$ bundle is the sum of its projections on to the tautological and tangent bundles.
@@
 whose angular part is the metric of the round 3-sphere of radius $r$.
-Each of these 3-spheres is a $U(2)$ orbit.
-and there is a unique $U(2)$ fixed point at $r=0$.
+Each of these 3-spheres is a $U(2)$ orbit, and there is a unique $U(2)$ fixed point at $r=0$.
 Note that the ratio of the circle radius to the 2-sphere radius is a constant, independent of $r$.
@@
 For us, the important observation is that for small positive $r$, the radius ratio is close to unity, so the squashed 3-sphere is approximately round.
-The consequence for our model is that the custodial $SO(4)$ symmetry – the symmetry of a round 3-sphere – will be approximately realised by the Higgs potential.
+The consequence for our model is that the custodial $SO(4)$ symmetry – the symmetry of a round 3-sphere – will be approximately realised.
 As $r\to\infty$, the radius ratio approaches zero, compatibly with the completion of $\mathbb{CP}^{2}$ at infinity being simply a round 2-sphere, which is a special 2-dimensional orbit of $U(2)$.
-It has radius $\frac{1}{2}$ and is at a finite distance from any other point in $\mathbb{CP}^{2}$.
+This 2-sphere has radius $\frac{1}{2}$ and is at a finite distance from any other point in $\mathbb{CP}^{2}$.
 4 Higgs Kinetic Term
@@
 $\Phi=(M,\phi_{1},\phi_{2})=(M,\phi)\,,$ (9)
-where $\phi=(\phi_{1},\phi_{2})$ is a complex scalar doublet.333 All these multiplets, and the left-handed fermion doublets below, are column vectors.
+where $\phi=(\phi_{1},\phi_{2})$ is a complex scalar doublet.444 All these multiplets, and the left-handed fermion doublets below, are column vectors.
 The Higgs field has mass dimension $1$, so the positive constant $M$ is a mass parameter.
@@
 $\Phi\to U\Phi=(M,e^{i\alpha/2}u\phi)\,.$ (11)
-In our model, as in the Standard Model, it is this $U(2)$ group that is gauged.
+In our model, as in the SM, it is this $U(2)$ group that is gauged.
 We introduce the $SU(2)$ and $U(1)$ gauge potentials, respectively $A_{\mu}$ and $B_{\mu}$, with $A_{\mu}$ hermitian and $B_{\mu}$ real.
@@
 where the prefactor $M^{2}$ is required to obtain an expression of mass dimension $4$.
-This kinetic term has novel features compared with the Standard Model kinetic term $\overline{D_{\mu}\phi}D^{\mu}\phi$, but reduces to it if $|\phi|$ is small compared with $M$.
+This kinetic term has novel features compared with the SM kinetic term $\overline{D_{\mu}\phi}D^{\mu}\phi$, but reduces to it if $|\phi|$ is small compared with $M$.
 5 Higgs Potential
@@
 $V$ has a local maximum value $\frac{1}{4}\lambda v^{4}$ at the point $\phi=0$, and a further local maximum value $\lambda M^{4}$ on the 2-sphere at infinity.
-As usual, the constant, vacuum field configuration is selected to be $\phi=(\phi_{1},\phi_{2})=\left(0,\frac{1}{\sqrt{2}}v\right)$, which is charge-neutral and real.
-The unbroken gauge group is then the usual electromagnetic $U(1)_{\rm em}$ subgroup of $U(2)$ with typical element
+As usual, in unitary gauge, the vacuum field configuration is selected to be $\phi=(\phi_{1},\phi_{2})=\left(0,\frac{1}{\sqrt{2}}v\right)$, which is charge-neutral and real.
+The unbroken gauge group is then the electromagnetic $U(1)_{\rm em}$ subgroup of $U(2)$ with typical element
 $u_{\rm em}=\begin{pmatrix}e^{i\beta}&0\\ 0&1\end{pmatrix}\,.$ (18)
@@
 $F_{\mu\nu}$ $=$ $\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}-ig[A_{\mu},A_{\nu}]\,,$ (19)
-$B_{\mu\nu}$ $=$ $\partial_{\mu}B_{\nu}-\partial_{\mu}B_{\nu}\,.$ (20)
+$B_{\mu\nu}$ $=$ $\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}\,.$ (20)
 Using the expansion of the $SU(2)$ gauge potential, $A_{\mu}=A_{\mu}^{a}I_{a}$, the $SU(2)$ field tensor acquires the expansion $F_{\mu\nu}=F_{\mu\nu}^{a}I_{a}$, where
@@
 The projection operator $1-P$ will later be key to the construction of gauge-invariant Yukawa terms.
-These assumptions mean that the kinetic terms (involving gauge-invariant spacetime derivatives) for the fermions are exactly as in the Standard Model, making use of the Dirac matrices $\gamma^{\mu}$ and $\gamma^{5}$.
+These assumptions mean that the kinetic terms (involving gauge-covariant spacetime derivatives) for the fermions are exactly as in the Standard Model, making use of the Dirac matrices $\gamma^{\mu}$ and $\gamma^{5}$.
 Fermion fields are doublet 4-spinors $\psi_{L}$ and singlet 4-spinors $\psi_{R}$ satisfying, respectively, $\gamma^{5}\psi_{L}=-\psi_{L}$ and $\gamma^{5}\psi_{R}=\psi_{R}$.
@@
 For the first-generation lepton pair, the electron and its neutrino, the fields are the doublet $l_{L}=(\nu_{L},e_{L})$ and the one singlet $e_{R}$.
-For the second- and third-generation lepton pairs we again do not write the Lagrangians explicitly.
+For the similar second- and third-generation lepton pairs we again do not write the Lagrangians explicitly.
 The lepton kinetic terms are
@@
 Matrix multiplication by $u$ acts the same way on the left and right columns of $\phi_{m}$, and therefore $\phi$ and $\phi_{C}$ transform the same way under $SU(2)$.
-We now turn to our $\mathbb{CP}^{2}$ model.
+We now turn to the $\mathbb{CP}^{2}$ model.
 There is still no unification of $\phi$ and $\phi_{C}$, and two distinct types of Yukawa term are needed.
@@
 Note that the projector $1-P$ acting on a right-handed fermion singlet produces a doublet and a singlet.
-The doublet plays the role in our $\mathbb{CP}^{2}$ model that the product of $\phi$ with the right-handed singlet does in the Standard Model, and its contraction with a left-handed fermion doublet is $U(2)$ invariant.
+The doublet plays the role in the $\mathbb{CP}^{2}$ model that the product of $\phi$ with the right-handed singlet does in the Standard Model, and its contraction with the corresponding left-handed fermion doublet is $U(2)$ invariant.
+This is the term we use.
 One could generate further Yukawa terms by using the projected singlet, but these would be of higher order in $\phi$.
@@
 The right-handed u-quark singlet is the 3-vector $(u_{R},0,0)$.
-Acting with $1-P_{C}$ on this, and contracting with $q_{L}$, gives the Yukawa term
+Acting with $1-P_{C}$ on this, and again contracting with $q_{L}$, gives the Yukawa term
 $\frac{-\Lambda_{u}M^{2}}{M^{2}+|\phi_{1}|^{2}+|\phi_{2}|^{2}}(\overline{u_{L}}\,\phi_{2}^{*}\,u_{R}-\overline{d_{L}}\,\phi_{1}^{*}\,u_{R})\,.$ (44)
@@
 But there remain some significant modifications to SM relations, and we will use these together with recent experimental results as reviewed by the Particle Data Group (PDG) [7] to place constraints on the BSM mass parameter $M$.
-Also, when the denominator factors are expanded out, there arise several dimension-6 terms that are Standard Model terms multiplied by $\bar{\phi}\phi$.
-In the following, we will explore the consequences of the $\mathbb{CP}^{2}$ SMEFT for (i) gauge boson masses, (ii) Higgs cubic and quartic couplings, (iii) Yukawa couplings.
+Also, when the denominator factors are expanded out, there arise several dimension-6 terms that are SM terms multiplied by $\bar{\phi}\phi$.
+In the following, we will explore the consequences of the $\mathbb{CP}^{2}$ SMEFT for (i) gauge boson masses, (ii) Higgs cubic and quartic couplings, (iii) Yukawa couplings, and (iv) Higgs boson couplings to gauge bosons.
 Constraints on $M$ currently come mainly from the $W$-boson to $Z$-boson mass ratio.
@@
 $Z_{\mu}=\cos\theta_{w}\,A_{\mu}^{3}-\sin\theta_{w}\,B_{\mu}\,,\quad A_{\mu}=\sin\theta_{w}\,A_{\mu}^{3}+\cos\theta_{w}\,B_{\mu}\,.$ (51)
-The $W$-boson fields are $W_{\mu}^{\pm}=\frac{1}{\sqrt{2}}(A_{\mu}^{1}\mp A_{\mu}^{2})$.
+The $W$-boson fields are $W_{\mu}^{\pm}=\frac{1}{\sqrt{2}}(A_{\mu}^{1}\mp iA_{\mu}^{2})$.
 It is not surprising that both gauge-boson mass terms (49) have no contribution from the photon field.
@@
 The combined gauge-boson mass term, including the prefactors, is
-$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\biggl{[}\left(M^{2}+\frac{1}{2}v^{2}\right)\frac{v^{2}}{8}\left[g^{2}(A_{\mu}^{1})^{2}+g^{2}(A_{\mu}^{2})^{2}+(g^{2}+g^{\prime 2})Z_{\mu}^{2}\right]$
-$\qquad\qquad\quad\qquad\qquad\qquad-\frac{v^{4}}{16}(g^{2}+g^{\prime 2})Z_{\mu}^{2}\biggr{]}.$ (52)
-$v^{4}Z_{\mu}^{2}$ cancels here, and what remains can be reexpressed as
-$\frac{1}{2}\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\left(\frac{gv}{2}\right)^{2}\left[\left(M^{2}+\frac{1}{2}v^{2}\right)((A_{\mu}^{1})^{2}+(A_{\mu}^{2})^{2})+\frac{M^{2}}{\cos^{2}\theta_{w}}Z_{\mu}^{2}\right].$ (53)
+$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\biggl{[}\left(M^{2}+\frac{1}{2}v^{2}\right)\frac{v^{2}}{8}\left[2g^{2}W_{\mu}^{-}W^{\mu+}+(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\right]$
+$\qquad\qquad\quad\qquad\qquad\qquad-\frac{v^{4}}{16}(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\biggr{]}.$ (52)
+$v^{4}Z_{\mu}Z^{\mu}$ cancels here, and what remains can be reexpressed as
+$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\left(\frac{gv}{2}\right)^{2}\left[\left(M^{2}+\frac{1}{2}v^{2}\right)W_{\mu}^{-}W^{\mu+}+\frac{1}{2}\frac{M^{2}}{\cos^{2}\theta_{w}}Z_{\mu}Z^{\mu}\right].$ (53)
 From this we read off the gauge boson masses
 $m_{W}=\frac{M}{\sqrt{M^{2}+\frac{1}{2}v^{2}}}\left(\frac{gv}{2}\right)\,,\quad m_{Z}=\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)}\left(\frac{gv}{2}\right)\frac{1}{\cos\theta_{w}}\,.$ (54)
-From here on, we work to leading order in $\frac{v^{2}}{M^{2}}$, as our estimates below show that this quantity is less than $0.01$. The masses become
+From here on, we work to leading non-trivial order in $\frac{v^{2}}{M^{2}}$, as the estimates below show that this quantity is less than $0.01$. The masses become
 $m_{W}=\left(1-\frac{v^{2}}{4M^{2}}\right)\left(\frac{gv}{2}\right)\,,\quad m_{Z}=\left(1-\frac{v^{2}}{2M^{2}}\right)\left(\frac{gv}{2}\right)\frac{1}{\cos\theta_{w}}\,.$ (55)
@@
 9.2 Higgs-boson mass and couplings
-Here we consider the terms of the Lagrangian involving the Higgs field alone, leaving out couplings of the Higgs field to the gauge potentials and fermions.
+Here we consider the terms of the Lagrangian involving the Higgs field alone, leaving out couplings of the Higgs field to the gauge bosons and fermions.
 In the SM, these terms are
@@
 $\phi=\frac{1}{\sqrt{2}}(0,v_{\rm SM}+H)\,,$ (61)
-where $v_{\rm SM}$ is the SM Higgs vev and $H$ the real Higgs boson field. Expanding out, we obtain
+where $v_{\rm SM}$ is the SM Higgs vev and $H$ the physical Higgs boson field. Expanding out, we obtain
 ${\mathcal{L}}_{\rm Higgs}=\frac{1}{2}\partial_{\mu}H\partial^{\mu}H-\lambda v_{\rm SM}^{2}H^{2}-\lambda v_{\rm SM}H^{3}-\frac{1}{4}\lambda H^{4}\,.$ (62)
@@
 However, the discussion here has clarified the need for the rescaling of the Higgs boson field $h$ as well as the Higgs vev $v$.
-These rescalings have an important effect on the Yukawa couplings, which we calculate next.
+These rescalings have an important effect on the Yukawa couplings, which we calculate next, and on the couplings of the Higgs boson to the gauge bosons.
 9.3 Higgs-fermion couplings
@@
 The kinetic terms for fermions have their canonical form in ${\mathcal{L}}$, so we need only investigate the Yukawa terms.
-To determine the fermion masses and coupling we need to retain only the terms up to linear order in the Higgs field $H$.
+To determine the fermion masses and couplings we need to retain only the terms up to linear order in the Higgs field $h$.
 As before, we convert the denominator factor in ${\mathcal{L}}_{\rm Yukawa}$ to a numerator factor and obtain for any fermion $f$,
@@
 $=$ $-\frac{1}{\sqrt{2}}\Lambda_{f}\overline{f}f\left(\left(1-\frac{v^{2}}{2M^{2}}\right)v+\left(1-\frac{3v^{2}}{2M^{2}}\right)h+O(h^{2})\right).$
-We next rescale $v$ and $h$ by the common factor $\left(1-\frac{v^{2}}{2M^{2}}\right)$ to obtain the normalised quantities $v_{\rm SM}$ and $H$. To linear order in $H$, the Yukawa term becomes
+We next rescale $v$ and $h$ by the common factor $1-\frac{v^{2}}{2M^{2}}$ to obtain the canonical $v_{\rm SM}$ and $H$. To linear order in $H$, the Yukawa term becomes
 ${\mathcal{L}}_{\rm Yukawa}=-\frac{1}{\sqrt{2}}\Lambda_{f}\overline{f}f\left(v_{\rm SM}+\left(1-\frac{v_{\rm SM}^{2}}{M^{2}}\right)H\right).$ (75)
@@
 Note that this lower bound does not conflict with the CDF result for the $W$-boson to $Z$-boson mass ratio.
+9.4 The $WWH$ and $ZZH$ couplings
+Finally, we determine the tree-level couplings of the $Z$- and $W$-bosons to the physical Higgs boson.
+These couplings are found by expanding the Higgs kinetic Lagrangian (14) to linear order in the Higgs field $h$.
+The coupling of the photon to the Higgs boson vanishes at tree level.
+As before, we go to unitary gauge, writing $\phi=\frac{1}{\sqrt{2}}(0,v+h)$, and will rescale both $v$ and the field $h$ to obtain the canonical SM mass for the $Z$-boson and the canonical kinetic term for the Higgs field.
+The rescaled quantities are
+$v_{\rm SM}=\left(1-\frac{v^{2}}{2M^{2}}\right)v\,,\quad H=\left(1-\frac{v^{2}}{2M^{2}}\right)h\,.$ (78)
+The desired coupling terms are quadratic in the gauge boson fields and linear in the Higgs field, and arise from (14) after excluding the derivative $\partial_{\mu}h$. We set
+$M^{2}+\bar{\phi}\phi=M^{2}+\frac{1}{2}(v+h)^{2}$ (79)
+and from (14) read off the remaining part
+$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}(v+h)^{2}\right)^{2}}\times$
+$\qquad\biggl{[}\left(M^{2}+\frac{1}{2}(v+h)^{2}\right)\frac{(v+h)^{2}}{8}\left[2g^{2}W_{\mu}^{-}W^{\mu+}+(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\right]$
+$\qquad\qquad\qquad-\frac{(v+h)^{4}}{16}(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\biggr{]},$ (80)
+where, similarly as before, the $(v+h)^{4}Z_{\mu}Z^{\mu}$ contributions cancel, leaving
+$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}(v+h)^{2}\right)^{2}}\biggl{[}M^{2}\frac{(v+h)^{2}}{8}\left[2g^{2}W_{\mu}^{-}W^{\mu+}+(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\right]$
+$\qquad\qquad\quad\qquad\qquad\qquad+\frac{(v+h)^{4}}{8}g^{2}W_{\mu}^{-}W^{\mu+}\biggr{]}.$ (81)
+Expanding in powers of $h$, we reproduce at zeroth order the masses of the gauge bosons, and at first order obtain the desired coupling terms
+$\frac{1}{2}\frac{M^{4}}{(M^{2}+\frac{1}{2}v^{2})^{2}}\,g^{2}v\,W_{\mu}^{-}W^{\mu+}h+\frac{1}{4}\frac{M^{4}(M^{2}-\frac{1}{2}v^{2})}{(M^{2}+\frac{1}{2}v^{2})^{3}}\,(g^{2}+g^{\prime 2})v\,Z_{\mu}Z^{\mu}h\,.$ (82)
+Now we convert to the rescaled quantities (78), and expand to linear order in $\frac{v_{\rm SM}^{2}}{M^{2}}$ to obtain
+$\frac{1}{2}g^{2}v_{\rm SM}\,W_{\mu}^{-}W^{\mu+}H+\frac{1}{4}(g^{2}+g^{\prime 2})\left(1-\frac{v_{\rm SM}^{2}}{M^{2}}\right)v_{\rm SM}\,Z_{\mu}Z^{\mu}H\,.$ (83)
+The couplings are therefore
+$g_{WWH}=\frac{1}{2}g^{2}v_{\rm SM}\,,\quad g_{ZZH}=\frac{1}{4}(g^{2}+g^{\prime 2})\left(1-\frac{v_{\rm SM}^{2}}{M^{2}}\right)v_{\rm SM}\,.$ (84)
+In the SM, these couplings are
+$g_{WWH}=\frac{1}{2}g^{2}v_{\rm SM}\,,\quad g_{ZZH}=\frac{1}{4}(g^{2}+g^{\prime 2})v_{\rm SM}\,,$ (85)
+so the $W$-boson coupling is unchanged at this order in the $\mathbb{CP}^{2}$ SMEFT, but the $Z$-boson coupling is reduced by the factor $1-\frac{v_{\rm SM}^{2}}{M^{2}}$.
+The current experimental constraints on these couplings from ATLAS and CMS at the Large Hadron Collider are quite tight [7].
+The measurements agree with the SM predictions to within 5%.
+If, for the sake of argument, one assumes the $Z$-boson coupling is less than 1% below the SM value, then $M$ would be constrained to be greater than 2.5 TeV.
 10 Conclusions
@@
 I am grateful to Mia West and Eetu Loisa for stimulating discussions about SMEFTs, initiated at the STFC HEP Forum: Completing the Higgs-saw puzzle, Cosener’s House, Abingdon, November ’23.
-I also thank Alex Mitov for comments.
-Additionally, I wish to thank the anonymous referee for several constructive suggestions which have been incorporated in this revised paper.
+I also thank Alex Mitov for comments, the anonymous referee for several constructive suggestions, and finally Alessandra Cappati for recommending calculation of the $WWH$ and $ZZH$ couplings during the IOP conference WIN 2025, Univ.
+Sussex, June ’25.
 This work has been partially supported by STFC consolidated grant ST/P000681/1.